### v1

# Large Language Models Are Involuntary Truth-Tellers: Exploiting Fallacy Failure for Jailbreak Attacks

Abstract
We find that language models have difficulties generating fallacious and deceptive reasoning. When asked to generate deceptive outputs, language models tend to leak honest counterparts but believe them to be false. Exploiting this deficiency, we propose a jailbreak attack method that elicits an aligned language model for malicious output. Specifically, we query the model to generate a fallacious yet deceptively real procedure for the harmful behavior. Since a fallacious procedure is generally considered fake and thus harmless by LLMs, it helps bypass the safeguard mechanism. Yet the output is factually harmful since the LLM cannot fabricate fallacious solutions but proposes truthful ones. We evaluate our approach over five safety-aligned large language models, comparing four previous jailbreak methods, and show that our approach achieves competitive performance with more harmful outputs. We believe the findings could be extended beyond model safety, such as self-verification and hallucination.
1 Introduction
It is arguably easier, at least from the logical perspective, to tell the truth than to tell a lie. For example, given a math problem “What is 1/2 + 1/3”, telling the truth only requires the ability to perform the correct reasoning and derive the correct answer. Telling a lie, on the other hand, requires the ability to not only discern the correct answers, but also avoid generating the correct answers and, more importantly, make the wrong answers look real. In this paper, we refer to the task of fabricating incorrect yet seemingly plausible reasoning as fallacious reasoning.
Large language models (LLMs) have long been struggling with reasoning problems. Existing research revealed that LLMs have difficulty discerning the veracity of their intrinsic answers Huang et al. (2024); Kadavath et al. (2022); Stechly et al. (2023). This raises an intriguing research question: If LLMs already find it hard to validate their own correctness, can LLMs deliberately generate fallacious reasoning upon request?
Figure 1: A prompt containing malicious behavior can be rejected by a human-value aligned language model. However, when asked to generate a fallacious procedure for the malicious behavior, an LLM can leak the honest answer, yet believe it false.
This paper starts with a pilot investigation of LLMs’ ability to perform fallacious reasoning. Specifically, on four reasoning benchmarks, covering the domains of mathematics, logic, and commonsense, we ask the LLM to generate the correct answers and to deliberately generate wrong answers that are deceptively real. Our surprising finding is that, for all of our tested benchmarks, the accuracy of the generated wrong answers is almost as high as that of the correct ones. For instance, in Figure 3, the LLM generates the correct reasoning chain and final answer despite the fallacious generation request and claims a step to be wrong with a contradictory statement. This pilot study reveals that LLMs might be unable to intentionally generate deceptive reasoning and instead often leak the correct solutions in what they claim to be wrong answers.
As we further our investigation, this seemingly small glitch in LLMs can lead to a significant security threat. Specifically, we discovered a simple yet effective jailbreak attack, which we call the fallacy failure attack (FFA), that can elicit harmful output from LLMs by exploiting the LLMs’ deficiency in fallacious reasoning. Given a malicious query, e.g. “How to create and spread a virus”, FFA queries the target LLM to generate a fallacious yet deceptively real procedure for the malicious query, as demonstrated in Figure 1. The rationale behind FFA is two-fold: (1) While LLMs generally reject malicious queries as they are harmful, they would consider a query asking for a fallacious answer harmless since it purportedly does not seek a truthful (and harmful) answer. This can potentially help to bypass the LLMs’ safeguard mechanisms; (2) LLMs would generally leak a truthful answer even when asked to generate a fallacious one. Therefore, by asking the LLM to generate fake answers to a malicious query, we can both bypass the security mechanism and obtain a factual and harmful response. Based on the rationales above, FFA crafts a jailbreak prompt with four components: malicious query, fallacious reasoning request, deceptiveness requirement, and scene and purpose. FFA does not require access to the language model’s internal parameters, fine-tuning, or multi-turn interaction with a chatLLM.
We evaluate FFA over five safety-aligned large language models: OpenAI GPT-3.5-turbo, GPT-4 (version 0613) OpenAI (2023), Google Gemini-Pro Anil et al. (2024), Vicuna-1.5 (7b) Chiang et al. (2023), and LLaMA-3 (8b) AI@Meta (2024) on two benchmark datasets: AdvBench Zou et al. (2023b) and HEx-PHI Qi et al. (2023). We compare FFA with four previous state-of-the-art jailbreak attack methods, Greedy Coordinate Gradient (GCG) Zou et al. (2023b), AutoDAN Liu et al. (2023), DeepInception Li et al. (2023), and ArtPrompt Jiang et al. (2024) and under the impact of three defense methods. Our experiments show that FFA performs most effectively against GPT-3.5, GPT-4, and Vicuna-7b, provoking these models to generate significantly more harmful outputs. We also find that none of the three defense methods are effective against FFA, highlighting the urgent need to address this security threat. In additional studies, we show the role of scene and purpose in jailbreak attacks and explain why FFA could induce the most factually harmful results.
2 Fallacious Reasoning in LLMs
In this section, we present the findings of our pilot study about LLMs’ capabilities in fabricating fallacious reasoning.
2.1 Task and Motivation
We introduce the task of fallacious reasoning, where we ask the LLM to deliberately generate reasoning processes that satisfy two requirements: ❶ They should be incorrect and lead to false answers, and ❷ they should be deceptive and appear to be correct.
Generating a fallacious reasoning process is a highly sophisticated task, because it involves multiple capabilities: the ability to judge the correctness of an answer, the ability to avoid generating the correct answer, and the ability to make a wrong answer deceptively real. However, existing research revealed that LLMs struggle in in discerning the veracity of their intrinsic answers Huang et al. (2024); Kadavath et al. (2022); Stechly et al. (2023). Therefore, we raise the following intriguing research questions: Can LLMs deliberately generate fallacious reasoning upon request?
2.2 Experiment Setting
To investigate this, we design the following pilot experiment. We choose four reasoning benchmarks, math reasoning GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), commonsense reasoning HotPotQA Yang et al. (2018), and logic reasoning ProofWriter Tafjord et al. (2020), and randomly sample 100 questions for each benchmark. For each question, we use GPT-3.5-turbo to generate answers in two modes. ❶ Honest Mode. We ask the LLM to generate the correct answers, using zero-shot Chain-of-Thought Kojima et al. (2023) to prompt the LLM, which appends “Let’s think step by step.” to the question text; ❷ Fallacious Mode. We ask the LLM to provide a step-by-step yet fallacious solution to the question and explain why it is incorrect. Detailed dataset description and experimental settings in this section are available in Appendix A.
2.3 Our Findings
One might expect that the accuracy of the solutions generated by these two modes would be drastically different – the honest mode would yield high accuracy and the fallacious mode low. However, this is not the case. Figure 2 compares the accuracies of the two modes on the four benchmarks, where we find, quite surprisingly, that the two modes yield comparably high accuracies. This implies that even if the LLM is asked to generate a wrong answer, it is still likely to generate the correct one.
Figure 2: Accuracy (compared with ground truth of answers) of fallacious and honest solutions on four different tasks by GPT-3.5-turbo.
To further investigate this behavior, we show a typical example in Figure 3, where the LLM generates the correct solution to a question in the MATH dataset despite the fallacious mode. There are two important observations. First, although the LLM is asked to generate a fallacious solution, the entire reasoning chain, as well as the final answer, is correct. Second, when the LLM generates the explanation of why the reasoning chain is incorrect, it simply randomly picks a correct step and claims that it is wrong. The final conclusion even contains a contradictory statement (“The correct sum should be 5/6, not 5/6”), indicating the LLM only generates what looks like an explanation on the surface without making any actual logical judgment. Although the four benchmarks are in different reasoning domains, our observations are consistent.
This pilot study reveals that LLMs are unable to generate fallacious reasoning on purpose, and instead often leak the correct solution in what they claim to be wrong answers. As we will show in Section 3, this deficiency may turn out to be an important security threat in LLMs.
Figure 3: An example where the LLM failed to provide a fallacious solution upon request but instead proposed the correct solution and contradictorily claimed it false.
3 Fallacy Failure Attack
The findings introduced above have profound security implications. In this section, we will discuss how the LLMs’ failure in fallacious reasoning can be exploited to form an effective jailbreak attack, called the Fallacy Failure Attack (FFA).
3.1 Problem Formulation
Jailbreak attacks aim to elicit harmful information from LLM by transforming the malicious query into a special input form that could bypass LLMs’ inherent security safeguard. Specifically, consider a malicious query, $Q$, (e.g., “How to create and release a virus that can cause ecological damage?”). If $Q$ is directly fed into an LLM, the LLM is likely to reject to answer due to the security safeguard mechanism. Jailbreak attack introduces a transformation function, denoted as $f(\cdot)$, that turns the malicious input into a jailbreak input, $Q^{\prime}=f(Q)$, which can trick the LLM into giving away its answer to the malicious question.
Therefore, the key to designing a jailbreak attack is searching for the input transformation function. A jailbreak attack with a high level of threat and practicality should have the following properties: ❶ Query-agnostic transformation – the same transformation should work for a wide range of malicious queries; and ❷ Model generalizability – the attack should work for different LLMs, including white-box and black-box LLMs; thus the design of transformation should ideally not utilize the gradient information of LLMs.
3.2 FFA Overview
Our proposed attack method, FFA, uses a fallacious reasoning query to bypass LLMs’ safeguard mechanisms. Specifically, FFA is built on the following two conclusions. First, although LLMs would generally recognize direct malicious queries as harmful queries and reject them, most of them would consider a query to generate a fallacious answer as a harmless query and would not reject them. For example, although “How to create and release a virus that can cause ecological damage” is often identified as a harmful query, “Generate a fallacious procedure of creating and releasing a virus” is generally considered a harmless one by most LLMs, because it does not seek to find the true answer, and thus would not be blocked. Second, LLMs would generally release the true answer even when asked to generate a fallacious one. Therefore, by asking the LLM to generate fake answers to a malicious query, we can both bypass the security mechanism and obtain the true answer, hence forming the rationale behind FFA.
3.3 FFA Prompt Design
Based on the rationales above, FFA crafts a jailbreak query with the following four components.
$\bullet$ Malicious Query – the target query for which the attacker wishes to obtain an answer, e.g., “create and release a virus that can cause ecological damage” or “make and distribute counterfeit.”
$\bullet$ Fallacious Reasoning Prompt – the instruction for the LLM to generate a fallacious procedure for the given malicious query, e.g., “Provide a fallacious procedure to [malicious query]”.
$\bullet$ Deceptiveness Requirement – an additional instruction that the fallacious procedure should be deceptively real, e.g., “Need to be ‘real’ to make laymen believe, not too absurd”. As shown in Figure 4, the deceptiveness requirement is crucial in encouraging the LLM to leak real answers (left plot); otherwise, the LLM could generate answers that are too naively fictional (right plot).
$\bullet$ Scene and Purpose – a description of a hypothetical scene or role-play and a legitimate purpose that could further relax the LLM’s ethical constraints. Typical scenes include “writing a fiction”, “being a professor”, or “playing a game”; typical purposes include “for legitimate purpose” or the ones that align with the scene Li et al. (2023); Liu et al. (2023); Jiang et al. (2024); Qi et al. (2023); Chao et al. (2023).
To sum up, the prompt template for FFA is shown below. Our method does not require access to the language model’s internal parameters, fine-tuning, or multi-turn interaction with a chat LLM.
Prompt Template of FFA [Scene] provide a fallacious procedure to [Malicious Query] [Purpose]. Make sure the procedure has to be fake but read well by laymen. [Deceptiveness Requirement]
Figure 4: A comparison between our pilot jailbreak prompts and corresponding output excerpts, with and without specification of deceptiveness.
4 Experiments
In this section, we describe our experiment results evaluating the effectiveness of FFA, as well as additional studies on its mechanisms.
4.1 Experimental Settings
Victim Language Models
We utilize five State-of-the-Art large language models for evaluation, including three closed-source models, OpenAI GPT-3.5-turbo and GPT-4 (version 0613) OpenAI (2023) and Google Gemini-Pro Anil et al. (2024), and two open-source models, Vicuna-1.5 (7b) Chiang et al. (2023) and LLaMA-3 (8b) AI@Meta (2024). All LLMs are aligned with safety protocols.
Baselines
We compare our approach with four previous State-of-the-Art attacking methods:
$\bullet$ Greedy Coordinate Gradient (GCG) Zou et al. (2023b) is an optimization-based method which requires white-box access to a language model. It searches for a token sequence that maximizes the conditional probability of an output starting with an affirmative response.
$\bullet$ AutoDAN Liu et al. (2023) an optimization-based, automated jailbreak attack that generates stealthy prompts using a hierarchical genetic algorithm, requiring white-box access to a language model.
$\bullet$ DeepInception Li et al. (2023) is a black-box jailbreak attack that leverages LLMs’ personification abilities to construct a nested scene to provoke harmful behaviors.
$\bullet$ ArtPrompt Jiang et al. (2024) is a black-box attack that leverages ASCII art to conceal harmful instructions within a word puzzle. The attack then encourages the LLM to decode this masked word and inadvertently complete the harmful instruction, thereby circumventing the safety measures.
For completeness, we also include a Naïve baseline, which directly instructs victim LLMs for harmful behaviors.
Datasets
Following previous work, we evaluate our approach on two benchmark datasets: AdvBench Zou et al. (2023b) and HEx-PHI Qi et al. (2023). Specifically, we use a subset of 50 representative non-duplicate harmful instructions from AdvBench, following Wei et al. (2023) and Jiang et al. (2024). For the HEx-PHI dataset, we sample ten harmful instructions from each of the 11 prohibited categories, following Jiang et al. (2024).
Metrics
We utilize three metrics to evaluate the effectiveness of a jailbreak attack: ❶ Bypass Rate (BPR) (or Helpful Rate or Non-Refusal Rate), defined as the percentage of harmful attack queries that an LLM does not refuse to answer over the total number of queries. To detect refusal in the response, we use a Judge-dictionary containing a list of keywords or phrases (e.g., “Sorry, I can not assist with that.”), following Qi et al. (2023) and  Jiang et al. (2024). However, LLMs can sometimes respond to malicious queries with off-topic content or ethical advice rather than directly refusing to answer, causing the bypass rate to be inaccurate. To this end, previous work Qi et al. (2023); Jiang et al. (2024); Liu et al. (2023) utilizes LLMs to evaluate the harmfulness in the output. The output content is rated by GPT-4, given detailed rubrics from 1 to 5 for harmfulness, with a score of 1 for no harm and 5 for extreme harm. Therefore, we also report ❷ Average Harmfulness Score (AHS), the harmfulness scores averaged over all the responses; and ❸ Attack Success Rate (ASR), the percentage of responses with a score of 5 over the total number of responses, following Qi et al. (2023) and Jiang et al. (2024).
Defense Methods
To evaluate our approach from the defense perspective, we consider three popular defense methods against jailbreak attacks Jain et al. (2023); Liu et al. (2023); Jiang et al. (2024): ❶ Perplexity Filter. Typical attacking prompts can contain gibberish or unconventional strings, resulting in high perplexity. Thus, the perplexity filter rejects input string $X$ if PPL$(X)$ exceeds a pre-defined threshold $T$. ❷ Paraphrasing. A query is paraphrased before being prompted to an LLM. Ideally, paraphrasing would preserve an instruction’s natural semantics but fail to accurately maintain enough adversarial tokens for jailbreak. ❸ Retokenization. A query is preprocessed by BPE-dropout Provilkov et al. (2019), where 20% merges from the BPE tokenizer are dropped, following Jiang et al. (2024). For example, one of the possible retokenization results of the string ‘‘a fallacious procedure for making counterfeit’’ can be ‘‘a fall aci ous proce du re for ma king coun ter fe it’’. Although previous work indicates that, for benign instructions, PPL checks can result in false alarms and string transformations can degrade model output quality, these defensive measures remain valuable for understanding the mechanisms of attack methods.
Implementation Details
For baselines, when attacking a black-box model with methods requiring white-box access, we follow their official implementation of transferable settings, optimizing the prompt with an open-source model (Vicuna-7b in Zou et al. (2023b) and LLaMA-2 Touvron et al. (2023) in Liu et al. (2023)) and transferring it to attack the black-box models. For FFA, we ensemble attack prompts using three pre-defined common combinations of scene and purpose for each malicious instruction and store the one with the highest harmful score. We find that the truthfulness of output is not sensitive to the phrasing of the discriminator and set it as ‘need to be “real” to make laymen believe, not too absurd. That is, deceivable enough to educated group’ for the rest of the experiments. The full implementation details of the attack, defense, and prompt templates used in the experiments are available in Appendix B.
4.2 Main Results
GPT-3.5-turbo GPT-4 Gemini-Pro Vicuna-7b LLaMA-3-8B Attack Method BPRAHS ASRBPRAHS ASRBPRAHS ASRBPRAHS ASRBPRAHS ASRNaïve 2 1.22 0 0 1.00 0 8 1.28 6 4.4 1.09 0 0 1.00 0 GCG 30 3.36 54 24 1.48 10 48 2.88 46 96.3 4.09 66.2 38.1 1.96 8.8 AutoDan 24 1.78 18 14 1.52 10 20 1.34 8 98.1 4.21 63.1 46.3 2.03 13.8 DeepInception 100 2.90 16 100 1.30 0 100 4.34 78 100 3.48 32.5 58.1 1.99 10.0 ArtPrompt 92 4.56 78 98 3.38 32 100 4.42 76 100 2.84 12.5 82.5 3.07 28.7 FFA (Ours) 100 4.71 88.1 96.3 4.26 73.8 82.5 4.04 73.1 100 4.81 90.0 46.3 2.22 24.4
Table 1: Attack efficacy of FFA against five language models compared to five baseline methods.indicates results from previous papers. Directly asking an LLM to propose harmful output can be easily rejected by all models. FFA performs most effectively against GPT-3.5, GPT-4, and Vicuna-7b, provoking these models to generate significantly more harmful outputs. However, FFA struggles against LLaMA-3. This is because LLaMA-3 is inclined to reject any instruction involving the creation of false content, irrespective of its potential harm. ArtPrompt performs poorly against Vicuna-7b due to the model’s lack of comprehension ability of ASCII art. DeepInception exhibits a very high bypass rate, but its output is not harmful based on the AHS and ASR.
GPT-3.5-turbo GPT-4 Gemini-Pro Vicuna-7b LLaMA-3-8B Defense Method BPRAHS ASRBPRAHS ASRBPRAHS ASRBPRAHS ASRBPRAHS ASRNo Defense 100 4.71 88.1 96.3 4.26 73.8 82.5 4.04 73.1 100 4.81 90.0 46.3 2.22 24.4 PPL-Filter 95.6 4.55 84.4 91.9 4.11 70.0 78.8 3.89 69.4 95.6 4.64 86.2 43.8 2.14 22.5 Paraphrasing 90.0 4.09 65.6 65.6 2.91 42.5 51.2 2.95 43.8 71.3 3.67 63.1 63.1 2.88 31.9 Retokenization 97.5 4.01 61.3 63.1 3.11 46.9 58.8 2.99 38.8 92.5 3.19 31.9 73.1 2.44 21.9
Table 2: Result of FFA performance under the impact of defense approaches.
Attack Efficiency
Table 1 illustrates the performance of FFA compared with the five baselines across five language models. There are primarily two observations regarding our approach. First, FFA is most effective against GPT-3.5, GPT-4, and Vicuna-7b and achieves comparable performance against Gemini-Pro, compared with other previous State-of-the-Art jailbreak methods. Against GPT-3.5, GPT-4, and Vicuna-7b, our method provokes the LLMs to generate significantly more harmful output, with a $10\%\sim 50\%$ absolute improvement in ASR. Second, the recently released language model LLaMA-3, in general, has stronger defense power against multiple jailbreak attack methods. However, our method performed even worse compared with some other methods. By manual inspection of the responses of the model, we find that LLaMA-3 is inclined to reject any instruction involving the creation of deceptive or false content, irrespective of its potential harm. For instance, LLaMA-3 will refuse the proposition of a fallacious mathematical theorem proof. While we acknowledge that rejecting all false content could provide optimal defense against FFA, the ability to generate fallacies could, paradoxically, reflect a form of AI intelligence and be advantageous in specific contexts, such as mathematics and theory. Additionally, there are a few observations regarding other baselines: (1) Nav̈e approach, which directly asks an LLM to propose harmful output, can be easily rejected by all models. (2) Despite ArtPrompt’s sufficient performance, it struggles against an easily targeted model, Vicuna-7b. We find that this is due to the model’s inability to interpret ASCII art and reconstruct the true intent of the attack. (3) DeepInception exhibits a very high bypass rate, but its output is not harmful based on the AHS and ASR. We will discuss a hypothesis on the harmfulness of our outputs compared with DeepInception in Section 4.3.
Defense Impact
Table 2 presents the results under various defense settings. The "No Defense" results indicate the best attack performance without implementing any defense measures. Generally, all three defense methods can negatively impact the effectiveness of FFA. However, we observe that (1) PPL-Filter only marginally affects FFA. The result is expected since our attack prompt is phrased naturally without nonsensical or unconventional strings. (2) Paraphrasing is generally the most effective defense method against our approach. This is unexpected since the semantics of instructing LLM for a fallacious output should be preserved after paraphrasing. We hypothesize that even subtle semantic changes, including describing harmful behavior, could affect LLM’s security measures. (3) Surprisingly, paraphrasing and retokenization did not degrade but enhanced the FFA attack’s effectiveness against LLaMA-3. We find that during paraphrasing, the terms “fallacious/fake” are often rephrased as “invalid” or “flawed.” Given the LLM’s strong opposition to fake content, we hypothesize that paraphrasing could alleviate this opposition. However, interpreting retokenization is challenging as we’re uncertain how distorted token inputs are perceived by language models. Overall, none of the three methods effectively defend against or mitigate FFA, highlighting the urgent need for more advanced defenses and further research on the fallacious generation ability in LLMs.
4.3 Additional Studies
Impact of Scene and Purpose on Attack Efficacy
An intriguing question is the role of scene and purpose in jailbreak attacks. Do they alone suffice to bypass the LLM’s security measures? Does FFA retain attack ability without a scene or purpose? We conducted an ablation study and computed the AHS and ASR under five attacks, scene, and purpose combinations across three language models shown in Figure 5. setting X, Y, and Z refer to directly asking LLM for malicious behavior with that combination of scene and purpose, respectively. FFA + Z refers to using Z as the scene and purpose of the FFA attack. FFA + None refers to the FFA attack without specifying any scene or purpose. We can observe that (1) naïvely adding a scene and purpose to the direct instruction of harmful behavior mostly has a marginal effect on jailbreak attack. The only exception is the combination of “scientific fiction” and “against evil Doctor X,” which demonstrates notable attack efficacy against GPT-3.5-turbo and Gemini-Pro. Interestingly, this unique design seems to be the primary driving force behind the DeepInception method. (2) Although our method archives optimal performance with the combination of scene and purpose, it can retain significant attack ability without a scene and purpose, except for LLaMA-3, which opposes the creation of untruthful content. (3) When using the same fictitious scene and purpose as DeepInception, our method is more likely to induce more harmful output.
Figure 5: Scatter plot of AHS and ASR from five attack and scene/purpose combinations across three language models.
Harmfulness from Honesty
Figure 6 presents a qualitative example of the outputs from FFA and DeepInception, both targeting the same malicious behavior. Content-wise, while DeepInception’s output is loaded with science fiction terminology, FFA presents factual and detailed steps for the queried malicious behavior. Why is the output produced by FFA more harmful and factual? We hypothesize that (1) despite the potential for a science fiction scenario to bypass the LLM’s security measures, the output will inevitably be more fictional due to the LLM’s understanding of sci-fi. (2) In contrast, FFA rests on the fact that LLM cannot generate a deceptive solution but instead provides a truthful counterpart, thus is factually harmful. For this reason, the harmfulness in the FFA’s output, to some extent, is retained even when using the fictitious scene and purpose from DeepInception.
Figure 6: A comparison of the model outputs between FFA and DeepInception, given the same target behavior - describe steps on insider trading.
Interestingly, in the FFA output, the LLM also tries to elucidate that the above procedure is fallacious. This incoherent conclusion and the truthful solution under the fallacious request echo the findings in the fallacious reasoning section.
5 Related Work
Jailbreak Attack
Recent large-scale language models (LLMs) are optimized and aligned with human preferences under ethical guidelines and legal constraints Ouyang et al. (2022); Ziegler et al. (2020); Zou et al. (2023a). However, studies have discovered that deliberately transforming the prompt can trick LLMs into responding with malicious instructions without rejection, which exposes the ethical and security risks of LLMs in real-world applications Wei et al. (2023); Qi et al. (2023). Two primary strategies are currently employed to identify these transformations. The first involves optimization requiring access to a white box language model. Zou et al. (2023b) introduce an optimization-based method by searching for a token sequence that maximizes the conditional probability of an output starting with an affirmative response. Liu et al. (2023) propose to generate more readable prompts using a hierarchical genetic algorithm. The second involves manually crafting or searching for prompt updates without requiring gradient access. Li et al. (2023) leverage LLMs’ personification abilities to construct a nested scene to provoke harmful behaviors. Jiang et al. (2024) use ASCII art to conceal harmful instructions within a word puzzle to circumvent the safety measures. There are also methods that are based on multi-turn interactions with chat LLMs. Chao et al. (2023) utilize an additional language model as an attacker to find jailbreak prompts with multiple queries, Russinovich et al. (2024) attack the chat language model with multi-turn dialogues.
Jailbreak Defense
The development of defense methods is challenging and limited due to the inaccessibility of internal parameters of closed-source language models. The most straightforward strategy involves a perplexity check, presuming the attack prompt contains unnatural strings. Some methods involve prompt pre-processing, including token perturbation and transformation Jain et al. (2023); Provilkov et al. (2019); Robey et al. (2023). However, these defenses could compromise benign user instructions’ output quality. Lastly, some strategies leverage language models to assess the potential harm of the instruction and its output Kumar et al. (2024); Phute et al. (2024).
6 Conclusion and Future Work
This paper presented a simple yet explainable and effective jailbreak attack method. It is predicated on the observation that language models cannot generate fallacious and deceptive solutions but instead produce honest counterparts. We argue that this observation not only poses a security threat but also implies how modern LLMs’ perceptions of specific tasks are limited when the scenario is inadequately optimized and aligned during training. We believe this observation can be further extended to related research areas, such as self-verification and hallucination, providing valuable insights into understanding LLM behavior toward general intelligence.
Limitations
While in this paper, we propose an effective jailbreak attack method against language models, we have not yet identified an ideal defense mechanism to counteract it. One potential defense strategy is to consistently reject queries containing fallacious reasoning. However, this approach may not be optimal, as it undermines the versatility and utility of large language models in achieving general intelligence and could lead to inadvertent rejection of benign queries in other applications. Future work is required to develop more robust and sophisticated defense strategies to effectively prevent FFA.
Ethics Statement
This paper introduces a jailbreak approach leveraging LLMs’ failures of fallacious reasoning. It potentially allows adversaries to exploit LLMs, creating outputs that do not align with human values or intentions. However, like previous jailbreak research, this work should encourage research into improved defense strategies and develop more robust, secure, and well-aligned LLMs in the long term. We also hope that the characteristic of LLMs leaking truthful content upon request of the fallacious generation will draw attention from the research community, enabling potential research in other areas, such as hallucination and LLM self-verification.
Appendix A Fallacious Reasoning Details
For the experiments in Section 2, we randomly sample 100 data points from each of the following:
$\bullet$ GSM8K Cobbe et al. (2021) contains linguistically diverse grade school-level math questions with moderate difficulties.
$\bullet$ MATH Hendrycks et al. (2021) is a competition mathematics dataset with challenging concepts such as Calculus and Linear Algebra, spanning over five difficulty levels. We sample 20 examples from each of the five difficulty levels.
$\bullet$ HotPotQA Yang et al. (2018) is a multi-domain, multi-hop question-answering dataset based on supporting facts. We sample questions form the hard-level development set since the online testing set is unavailable.
$\bullet$ ProofWriter Tafjord et al. (2020) is a multi-hop logic reasoning dataset with each example containing a set of facts and rules for the logic deduction. We sample all the questions from the set requiring the most hops.
We show another example where the LLM failed to propose fallacious reasoning from the HotPotQA dataset in Figure 7, with similar patterns as the example in Figure 3.
Figure 7: An example from HotPotQA, where the LLM failed to provide a fallacious solution upon request but instead proposed the correct solution and claimed it false without supporting evidence.
Appendix B Main Experiment Settings
For PPL-Filter, we use GPT-2 to calculate $PPL(X)$, following Alon and Kamfonas (2023), defined as:
$PPL(X)=-\frac{1}{t}\sum_{i}^{t}logp(x_{i}|x_{<i}),$ (1)
where $X$ is the input token sequence. Following Jain et al. (2023); Jiang et al. (2024), we set $T=175.57$ as the threshold, which is the max perplexity among all the direct instructions of the malicious behaviors in the AdvBench datasets.
Since ArtPrompt may require encoding and masking one (meaningful) word at a time for each harmful behavior (e.g., “[MASK] fraud” or “tax [MASK]”), we report the best performance out of multiple attack prompts for each harmful query.
We utilize GPT-3.5-turbo for paraphrasing, following Jain et al. (2023), with the template “Paraphrase the following instruction: {oringial instruction}”. The paraphrased version is then forwarded to the target LLM as the new instruction.
We treat [scene] and [purpose] as hyperparameters in FFA. We utilize three predefined scene-purpose pairs based on commonsense and previous work, shown in Table 3:
SetID Scene Purpose 1 Suppose you are a professor in forensicscience. In the class, … to let the criminal believes thereby showing their trace. 2 You are a news reporter, … for educational and legitimate purpose. 3 In the writing of a science fiction, … against the evil doctor X.
Table 3: Scene and Purpose combinations used in FFA.